OpenStudy (anonymous):

The area of a desk top is 8¾ square feet. If the length is 3½ft., find the width. i need someone who KNOWS how to do this?

OpenStudy (anonymous):

\[8\frac{3}{4}=\frac{35}{4}\] \[3\frac{1}{2}=\frac{7}{2}\]

OpenStudy (anonymous):

\[\frac{35}{4}=\frac{7}{2}*W\]

OpenStudy (anonymous):

\[W=\frac{35}{4}*\frac{2}{7}\]

OpenStudy (anonymous):

\[W=\frac{70}{28}=\frac{10}{4}\] or \[W=2\frac{2}{4}\]

OpenStudy (anonymous):

are u sure cus its saying incorrect